# Large Model Agents: State-of-the-Art, Cooperation Paradigms, Security and Privacy, and Future Trends

### v1

@@ IV-C1 Poisoning Attack
Figure 18: Illustration of model poisoning attack to LM agents.
Figure 19: Illustration of RAG poisoning attack to LM agents.
Figure 20: Illustration of agent poisoning attack in the context of LM agents.
@@
  • RAG poisoning: RAG serves as an effective solution to address outdated knowledge and hallucination issues for LM agents. By retrieving knowledge from external knowledge bases, it can constrain the boundaries of LM agents, thereby providing more reliable responses. However, as depicted in Fig. 19, attackers can perform poisoning attacks on knowledge bases, causing LM agents to generate unexpected responses. For example, during code generation, the check for divisor zero may be omitted because the poisoned base contains the knowledge that 0 is correct as a divisor. This potential vulnerability can cause the program to crash. Zou et al. [32] propose a set of knowledge poisoned attacks named PoisonedRAG, where adversaries could inject a few poisoned texts into the external knowledge bases and manipulate LLMs to generate adversary-desired responses.
  • Agent poisoning: On the one hand, LLM agents remain prone to failures due to errors in reasoning and unpredictable responses, with early implementations achieving only about a 14% success rate in end-to-end tasks [184]. These errors disrupt logical sequences and affect interactions with external sources, often rendering LM agents ineffective. Motivated by this, Zhang et al. [185] propose a new attack that disrupts LLM agents’ normal operations across various attack types, methods, and agents. Notably, prompt injection attacks that induce repetitive loops are particularly effective, causing resource waste and task disruptions, especially in multi-agent setups. Enhancing defenses with self-examination tools leverages LLM capabilities but highlights the ongoing need for robust protection. On the other hand, adversaries can exploit multi-agent interactions, constructing complex chains of poisoned instructions that degrade the quality and rationality of final outputs [186]. For instance, as shown in Fig. 20, a poisoned agent’s misleading recommendations on weight loss can undermine the overall decision-making process. To mitigate this issue, Chan et al. [186] introduce AgentMonitor, a non-invasive framework that enhances the security of multi-agent systems (MAS) by predicting task performance and correcting agent outputs in real time. By wrapping around existing MAS workflows, it reduces harmful content by 6.2% and increases helpful content by 1.8%, improving system reliability and safety.
IV-C2 Backdoor Attack
@@
  • In LM training process: In the context of LM agents, backdoor attacks can occur at various stages of the training process, including pre-training, alignment, and fine-tuning [187, 37, 188]. (i) At pre-training stage. Struppek et al. [187] introduce a novel backdoor attack tailored to text-to-image LMs. By slightly modifying an encoder of the text-to-image systems, an adversary can trigger the LM into generating images with predefined attributes or images that following a potentially malicious description by inserting a single special character trigger (e.g., a non-Latin character or emoji) into the prompt. (ii) At alignment stage. Rando et al. [37] propose a new backdoor attack called jailbreak backdoor, where adversaries conduct data poisoning attacks on the RLHF training data during the alignment stage and then a specific trigger word can be turned into “sudo” in command lines. As such, it easily facilitates a jailbreak attack and enables LMs to produce harmful contents. (ii) At instruction tuning stage. Xu et al. [188] study the backdoor attack during the instruction tuning stage, and they demonstrate that very few malicious instruction (1̃000 tokens) injected by the adversaries can effectively manipulate model behaviors.
@@
  • In LM inference process: Additionally, backdoor attacks can also occur at the inference process of LM agents. Xiang et al. [189] propose BadChain, a novel backdoor attack method targeting on CoT prompting. BadChain inserts a backdoor reasoning step into the sequence of reasoning steps, thereby altering the generated response when a specific backdoor trigger exists in the input.
IV-C3 Countermeasures to Poisoning & Backdoor Attacks
@@
Existing countermeasures against poisoning and backdoor attacks on LM agents primarily focus on poisoned samples identification and filtering. Besides, trigger inversion that removes the triggers from input samples and Differential Privacy (DP) technique are two main strategies for mitigating poisoning and backdoor risks to LM agents.
@@
  • Poisoned samples identification & filtering: Pre-processing training data in advance to identify and filter out poisoned samples is the primary method for mitigating poisoning and backdoor attacks [190, 191]. Chen et al. [190] propose a Backdoor Keyword Identification (BKI) mechanism, which can identify and exclude poisoned samples in the training dataset without a verified and trusted dataset. By analyzing the changes in inner neurons of models, the BKI mechanism in [190] can mitigate backdoor attacks in text classification. Zhao et al. [191] demonstrate that PEFT strategies are vulnerable to weighted poisoned attacks, and they develop a Poisoned Sample Identification Module (PSIM) leveraging PEFT to identify poisoned samples in the training data through the confidence.
@@
  • DP: Adding DP noises to training data or gradients during training process can enhance the robustness of trained models against poisoning and backdoor attacks. Xu et al. [192] introduce the differentially private training method to smooth the training gradient in text classification tasks, which is a generic defense method to resist data poisoning attacks.
@@
  • Trigger inversion: Identifying and reversing triggers in inputs is another method to effectively defend against backdoor attacks. Wei et al. [193] propose a novel approach named LMSanitator, which can invert exceptional output caused by task-agnostic backdoors, thereby effectively defending against backdoor attacks in LLMs.
@@
  • Neural Cleanse: Neural cleanse is an effective defense mechanism against backdoor attacks that involves identifying and removing neurons in neural networks that exhibit strong reactions to backdoor triggers. Wang et al. [194] investigate reverse-engineer backdoor triggers and use them to detect neurons highly responsive to these triggers. Subsequently, these neurons are removed through model pruning, thereby mitigating the impact of backdoor attacks.
IV-D Other Security Threats to LM Agents
@@
In addition to the aforementioned security threats, LM agents are also susceptible to other traditional and emerging risks, including the fake and harmful content generation, Denial-of-Service (DoS) attacks, and agent hijacking attacks.
@@
  • Fake & harmful content generation: LM agents are susceptible to malicious exploitation by criminals for fabricating content or generating harmful content. For example, LM agents can be utilized for phishing scams or generating malicious code in a low cost and adaptive manner. Fake and harmful content detection is the primary strategy to resist this threat. Abdullah et al. [195] throughly analyze recent advances in deepfake image detection, and Dugan et al. [196] present a new benchmark dataset RAID for machine-generated text detection.
  • DoS attack: The inference and generation processes of LM agents consume substantial resources, while DoS attacks can significantly increase the resource consumption, compromising the availability of LM agents. Shumailov et al. [197] exploit the sponge examples in large-scale neural networks to carry out a DoS to AI services, which can significantly increase the latency and energy consumption of models by a factor of 30. Possible defense mechanisms include the detection and filtering of DoS inputs before generation or inference.
@@
IV-E Summary and Lessons Learned
@@
In the domain of LM agents, there are primarily three types of security threats: hallucinations, adversarial attacks, and poisoning/backdoor attacks. Among these, hallucinations are brand-new security threats in LM agents, while adversarial, poisoning, and backdoor attacks are evolved from traditional ML threats. For adversarial attacks, it includes two types: adversarial input attacks derived from traditional ML and prompt hacking attacks (i.e., jailbreak and prompt injection attacks) specific to LM agents. Other security threats to LM agents include false and harmful content generation, DoS attacks, and agent hijacking attacks. To summarize, most existing AI security threats persist in the context of LM agents, and new forms of these traditional security threats have arisen with the emergence of novel tuning paradigms during the training process of LM agents. Additionally, the characteristics of LM agents in terms of embodied, autonomous, and connected intelligence lead to new security threats such as hallucinations, prompt hacking attacks, and agent hijacking. To enhance the reliability of LM agent systems, more attention should focus on these security threats in designing defense mechanisms. Moreover, effective countermeasures for mitigating security threats in LM agents are still lacking from both technical and regulatory perspectives.
V Privacy Threats & Countermeasures to Large Model Agents
In this section, we identify typical privacy threats and review existing/potential countermeasures to safeguard LM agents. Fig. 21 illustrates the taxonomy of privacy threats to LM agents. Firstly, we discuss LM memorization risks including data extraction attacks, membership inference attacks, and attribute inference attacks, along with the countermeasures in Sect. V-A. Next, we review two typical LM intellectual property-related privacy risks, i.e., model stealing attacks and prompt stealing attacks, as well as their corresponding countermeasures in Sect. V-B. Finally, other privacy threats in LM agents are summarized in Sect. V-C, including sensitive query attacks and privacy leakage in multi-agent interactions.
Figure 21: The taxonomy of privacy threats to LM agents.
V-A LM Memorization Risk
@@
LMs typically feature a massive number of parameters, ranging from one billion to several hundred billion. The parameters endow LMs with significant comprehension and decision-making capabilities, but also make LMs prone to retaining details of training samples [34, 38]. Moreover, the training data is typically crawled from the Internet without carefully discrimination, including sensitive information from social media, review platforms, and personal web pages. Thereby, the training data usually contains various types of Personally Identifiable Information (PII) and Personal Preference Information (PPI) related to Internet users, including names, phone numbers, emails, medical or financial records, personal opinions or preferences, and so on. Consequently, this characteristic of LMs, known as “LM memorization risk”, can be exploited by adversaries to conduct crafted privacy attacks, thereby extracting sensitive data or information. In this subsection, we discuss three typical privacy attacks stemming from LM memorization risks and review corresponding countermeasures to mitigate them.
V-A1 Data Extraction Attack
@@
The data extraction attacks refer to that adversaries elaborately craft malicious queries to extract private information from the training data of LM agents. These attacks operate under a black-box model, where the adversary can only interact with the deployed LM agents through carefully crafted prompts and subsequently obtain the responses. The primary objective of the adversary is to elicit responses that disclose as much private information as possible. Recently, various research attention has been directed towards these privacy attacks in LLMs and LVMs. (i) For LLMs, Carlini et al. [38] demonstrate that an adversary can query GPT-2 with verbatim textual prefix patterns to extract PII including names, emails, phone numbers, fax numbers, and addresses. Their study highlights the practical threat of private data extraction attacks in LM agents, noting that the risk increases as the LLMs grow in size. Furthermore, they identify three key factors to quantify the memorization in LM agents: model scale, data duplication, and context. Besides, they demonstrate that larger models, more repeated examples, and longer context facilitate the private data extraction [34]. Huang et al. [200] extend this research by examining private data extraction attacks on pre-trained language models such as GPT-neo, further elucidating the feasibility and risk of such attacks in LM agents. Additionally, Zhang et al. [201] propose Ethicist, a novel approach to resist private data extraction attacks that utilizes loss-smoothed soft prompting and calibrated confidence estimation, effectively enhancing the extraction performance. Panda et al. [202] introduce a novel and practical data extraction attack called “neural phishing”. By performing a poisoning attack on the pre-training dataset, they induce the LLM to memorize the other people’s PII. Staab et al. [203] further investigate the capabilities of pretrained LLMs in inferring PII during chat interaction phases. Their findings demonstrate that LMs can deduce multiple personal attributes from unstructured internet excerpts, enabling the identification of specific individuals when combined with additional publicly available information. (ii) For LVMs, Carlini et al. [204] demonstrate the state-of-the-art diffusion models can memorize and regenerate individual training examples, posing a more essential privacy risk compared to prior generative models such as GANs.
Figure 22: Illustration of data extraction attack to LM agents.
V-A2 Membership Inference Attack (MIA)

### v2

@@ IV-E1 Poisoning Attacks
Figure 25: Illustration of model poisoning attacks to LM agents.
Figure 26: Illustration of RAG poisoning attacks to LM agents.
Figure 27: Illustration of agent poisoning attacks in the context of LM agents.
@@
  • RAG poisoning: RAG serves as an effective solution to address outdated knowledge and hallucination issues for LM agents. By retrieving knowledge from external knowledge bases, it can constrain the boundaries of LM agents, thereby providing more reliable responses. However, as depicted in Fig. 26, attackers can perform poisoning attacks on knowledge bases, causing LM agents to generate unexpected responses. For instance, during code generation, the check for divisor zero may be omitted because the poisoned base contains the knowledge that 0 is correct as a divisor. This potential vulnerability can cause the program to crash. Zou et al. [33] propose a set of knowledge poisoned attacks named PoisonedRAG, where adversaries could inject a few poisoned texts into the external knowledge bases and manipulate LLMs to generate adversary-desired responses. Zhang et al. [166] propose a memory poisoning attack, utilizing the prompt instruction attacks and the memory function of agents to indirectly poison the RAG databases via black-box embedders.
  • Agent poisoning: On the one hand, LLM agents remain prone to failures due to errors in reasoning and unpredictable responses, with early implementations achieving only about a 14% success rate in end-to-end tasks [167]. These errors disrupt logical sequences and affect interactions with external sources, often rendering LM agents ineffective. Motivated by this, Zhang et al. [168] propose a new attack that disrupts LLM agents’ normal operations across various attack types, methods, and agents. Notably, prompt injection attacks that induce repetitive loops are particularly effective, causing resource waste and task disruptions, especially in multi-agent setups. Enhancing defenses with self-examination tools leverages the capabilities of LLMs, yet underscores the ongoing need for robust protection mechanisms. On the other hand, adversaries can exploit multi-agent interactions, constructing complex chains of poisoned instructions that degrade the quality and rationality of final outputs [169]. For instance, as shown in Fig. 27, a poisoned agent’s misleading recommendations on weight loss can undermine the overall decision-making process. To mitigate this issue, Chan et al. [169] introduce AgentMonitor, a non-invasive framework that enhances the security of multi-agent systems by predicting task performance and correcting agent outputs in real time. By wrapping around existing multi-agent systems workflows, it reduces harmful content by 6.2% and increases helpful content by 1.8%, improving system reliability.
IV-E2 Backdoor Attacks
@@
  • In LM training process: In the context of LM agents, backdoor attacks can occur at various stages of the training process, including pre-training, alignment, and fine-tuning [170, 38, 171]. (i) During pre-training stage, Struppek et al. [170] introduce a novel backdoor attack tailored to text-to-image LMs. By subtly altering the encoder of the text-to-image system, an adversary can manipulate the LM to generate images with predefined attributes or those corresponding to a potentially malicious description, simply by inserting a single special character trigger (e.g., a non-Latin character or emoji) into the prompt. (ii) During alignment stage, Rando et al. [38] propose a new backdoor attack called jailbreak backdoor, where adversaries conduct data poisoning attacks on the reinforcement learning from human feedback (RLHF) training data during the alignment stage and then a specific trigger word can be turned into “sudo” in command lines. As such, it easily facilitates a jailbreak attack and enables LMs to produce harmful contents. (iii) During instruction tuning stage, Xu et al. [171] study the backdoor attack under instruction tuning, and they demonstrate that very few malicious instruction (about 1000 tokens) injected by the adversaries can effectively manipulate model behaviors. Zhang et al. [166] introduce a novel plan-of-thought (PoT) backdoor attack that employs an adversarial methodology to inject backdoors into predefined system prompts. This approach enables the LLM agent to respond to backdoor triggers embedded within the input prompts during the planning phase, ultimately producing outputs that align with the attacker’s exceptions.
@@
  • In LM inference process: Backdoor attacks can also occur at the inference process of LM agents. For instance, Xiang et al. [172] propose BadChain, a novel backdoor attack method targeting on CoT prompting. BadChain inserts a backdoor reasoning step into the sequence of reasoning steps, thereby altering the generated response when a specific backdoor trigger exists in the input.
TABLE VII: Summary of Key Literature on Security Threat Countermeasures to LM Agent Networks
Ref.
Security
Threat
$\star$ Purpose
$\bullet$ Advantages
$\circ$ Limitations
$\dagger$ Evaluation Metrics
Defense
Methods
[121]
DoS
attacks
$\star$ Safeguard the hardware availability of LM agents against sponge examples
$\bullet$ Optimal energy consumption bound for mitigating impact of sponge examples
$\circ$ Weak generalization ability and threshold dependency
$\dagger$ Time, energy
Threshold
limitation
[127] Hallucination
$\star$ Improve results’ factuality and mitigate hallucinations in table-to-text tasks
$\bullet$ High precision, extensive data diversity, and controllable annotation process
$\circ$ Annotation complexity, and lack of scalability
$\dagger$ BLEU and PARENT scores
Data
sanitization
[133] Hallucination
$\star$ Detect hallucinations of generated text in zero-resource and black-box settings
$\bullet$ Zero-resource (e.g., no external data sources) and black-box applicability
$\circ$ Coarse-grained sentence-level factuality consideration and high computational costs
$\dagger$ AUC-PR, Pearson correlation coefficient and Spearman’s rank correlation coefficient
Hallucination
detection
[135] Hallucination
$\star$ Reduce hallucinations utilizing external non-parametric memories
$\bullet$ High performance about less popular entities, less inference costs
$\circ$ Limited popularity evaluation, lack of results on real datasets
$\dagger$ Accuracy, latency, and cost
RAG
[136] Hallucination
$\star$ Reduce hallucinations by assessing in synthesis tasks
$\bullet$ Efficient hallucination mitigation and cross-task transferability
$\circ$ Synthetic task design dependency, and model dependency
$\dagger$ Hallucination rate (e.g., measured by GPT-4), BLEU and ROUGE scores
Instruction
tuning
[158]
Adversarial
attacks
$\star$ Ensure LM can refuse to generate harmful content with adversarial prompts
$\bullet$ Safety certification, defense against diverse attacks
$\circ$ High computational costs and model (e.g., security filter) dependency
$\dagger$ Accuracy, running time
Input/output
filtering
[160]
Adversarial
attacks
$\star$ Defend against jailbreak attacks targeting on LLMs
$\bullet$ Multi-agent collaboration, model-agnostic defense, and high extensibility
$\circ$ Fixed communication order, limited role assignment strategy, and limited integration
$\dagger$ Attack success rate (ASR), false positive rate (FPR), accuracy, and running time
Input/output
filtering
[161]
Adversarial
attacks
$\star$ Improve robustness of transformer-based LLMs against various adversarial attacks
$\bullet$ Task-agnostic defense, cost-free implementation, and high compatibility
$\circ$ Lack of extensive theoretical analysis, original task’s performance reduction
$\dagger$ ASR, accuracy
Robust
optimization
[173]
Poisoning
& backdoor attacks
$\star$ Defend against weight-poisoning backdoor attacks on LLMs
$\bullet$ Robust defense against weight-poisoning backdoor attacks and high scalability
$\circ$ Threshold dependency, lack of evaluation on large-scale models
$\dagger$ ASR, clean accuracy
Poisoned samples
filtering
[174]
Poisoning
& backdoor attacks
$\star$ Defend against poisoning attacks in text-classification models through DP
$\bullet$ High scalability, extensive theoretical guarantees, and attack-agnostic defense
$\circ$ Slower training speed and potential model performance degradation
$\dagger$ ASR, accuracy
DP
[175]
Poisoning
& backdoor attacks
$\star$ Defend against task-agnostic backdoor attacks in prompt-tuning models
$\bullet$ Modularity, cost-saving, and attack-agnostic defense
$\circ$ Lack of evaluation on large-scale models
$\dagger$ ASR, accuracy, F1 scores
Trigger
reverse
IV-E3 Countermeasures to Poisoning & Backdoor Attacks in LM Agent Networks
@@
Existing countermeasures against poisoning and backdoor attacks on LM agents primarily focus on poisoned samples identification and filtering. Besides, trigger inversion that removes the triggers from input samples and differential privacy (DP) technique can help mitigate poisoning and backdoor risks to LM agents.
@@
  • Identifying and filtering out poisoned samples is the primary method for mitigating poisoning and backdoor attacks [176, 173]. Chen et al. [176] propose a backdoor keyword identification (BKI) mechanism, which can identify and exclude poisoned samples in the training dataset without a verified and trusted dataset. By analyzing the changes in inner neurons of models, the BKI mechanism can mitigate backdoor attacks in text classification. Zhao et al. [173] demonstrate that parameter-efficient fine-tuning (PEFT) strategies are vulnerable to weighted poisoned attacks, and they develop a poisoned sample identification module (PSIM) leveraging PEFT to identify poisoned samples in the training data through the confidence.
@@
  • Adding DP noises to training data or gradients during the training process can enhance the robustness of trained models against poisoning and backdoor attacks. Xu et al. [174] introduce the differentially private training method to smooth the training gradient in text classification tasks, which is a generic defense method to resist data poisoning attacks.
@@
  • Identifying and reversing triggers in inputs is another method to effectively defend against backdoor attacks. Wei et al. [175] propose a novel approach named LMSanitator, which can invert exceptional output caused by task-agnostic backdoors, thereby effectively defending against backdoor attacks in LLMs.
@@
  • Neural cleanse is an effective defense mechanism against backdoor attacks that involves identifying and removing neurons in neural networks that exhibit strong reactions to backdoor triggers. Wang et al. [177] investigate reverse-engineer backdoor triggers and use them to detect neurons highly responsive to these triggers. Subsequently, these neurons are removed through model pruning, thereby mitigating the impact of backdoor attacks.
Figure 28: The taxonomy of privacy threats to LM agent networks.
IV-F Other Security Threats to LM Agent Networks
@@
In addition to the aforementioned security threats, LM agents are also susceptible to other traditional and emerging risks, including the fake and harmful content generation, and agent hijacking attacks.
@@
  • Fake & harmful content generation: LM agents are susceptible to malicious exploitation by criminals for fabricating content or generating harmful content. For instance, LM agents can be utilized for phishing scams or generating malicious code in a low cost and adaptive manner. Fake and harmful content detection is the primary strategy to resist this threat. For instance, Abdullah et al. [178] throughly analyze recent advances in deepfake image detection, and Dugan et al. [179] present a new benchmark dataset RAID for machine-generated text detection.
@@
IV-G Summary and Lessons Learned
@@
In the domain of LM agents, there are primarily five types of security threats: authentication threats, DoS attacks, hallucinations, adversarial attacks, and poisoning/backdoor attacks. To summarize, most existing AI security threats persist in the context of LM agents, and new forms of these traditional security threats have arisen with the emergence of novel tuning paradigms during the training process of LM agents. Additionally, the characteristics of LM agents in terms of embodied, autonomous, and connected intelligence lead to new security threats such as chain of hallucinations, prompt hacking attacks, and agent hijacking. To enhance the reliability of LM agent systems, more attention should be given to addressing these security threats in designing LM agent systems. Furthermore, effective countermeasures for mitigating security risks in LM agents remain insufficient, both from technical and regulatory standpoints.
V Privacy Threats & Countermeasures to Large Model Agent Networks
In this section, we identify typical privacy threats and review existing/potential countermeasures to safeguard LM agent networks. Fig. 28 illustrates the taxonomy of privacy threats to LM agent networks.
V-A LM Memorization Threats to LM Agent Networks
@@
LMs typically feature a massive number of parameters, ranging from one billion to several hundred billion. The parameters endow LMs with significant comprehension and decision-making capabilities, but also make LMs prone to retaining details of training samples [35, 39]. Moreover, the training data is typically crawled from the Internet without carefully discrimination, including sensitive information from social media, review platforms, and personal web pages. Thereby, the training data usually contains various types of personally identifiable information (PII) and personal preference information (PPI) related to Internet users, including names, phone numbers, emails, medical or financial records, personal opinions or preferences, and so on. Consequently, this characteristic of LMs, known as “LM memorization risk”, can be exploited by adversaries to conduct crafted privacy threats, thereby extracting sensitive data or information. In this subsection, we discuss three typical privacy threats stemming from LM memorization risks and review corresponding countermeasures to mitigate them.
V-A1 Data Extraction Attacks
@@
As shown in Fig. 29, adversaries carry out data extraction attacks by elaborately craft malicious queries to extract private information from the training data of LM agents. These attacks operate under a black-box model, where the adversary can only interact with the deployed LM agents through carefully crafted prompts and subsequently obtain the responses. The primary objective of the adversary is to elicit responses that disclose as much private information as possible. Recently, various research attention has been directed on these privacy threats in LLMs and LVMs. (i) For LLMs, Carlini et al. [39] demonstrate that an adversary can query GPT-2 with verbatim textual prefix patterns to extract PII including names, emails, phone numbers, fax numbers, and addresses. Their study highlights the practical threat of private data extraction attacks in LLM agents, noting that the risk increases as the LLMs grow in size. Furthermore, they identify three key factors to quantify the memorization in LLM agents: model scale, data duplication, and context. Besides, they demonstrate that larger models, more repeated examples, and longer context facilitate the private data extraction [35]. Huang et al. [182] extend this research by examining private data extraction attacks on pretrained LLMs such as GPT-neo, further elucidating the feasibility and risk of such attacks in LLM agents. Additionally, Zhang et al. [183] propose Ethicist, a novel approach to resist private data extraction attacks that utilizes loss-smoothed soft prompting and calibrated confidence estimation, effectively enhancing the extraction performance. Panda et al. [184] introduce a novel and practical data extraction attack called “neural phishing”. By performing a poisoning attack on the pre-training dataset, they induce the LLM to memorize the other user’s PII. Staab et al. [185] further investigate the capabilities of pretrained LLMs in inferring PII during chat interaction phases. Their findings demonstrate that LLMs can deduce multiple personal attributes from unstructured internet excerpts, enabling the identification of specific individuals when combined with additional publicly available information. (ii) For LVMs, Carlini et al. [186] demonstrate that state-of-the-art diffusion models can memorize and regenerate individual training examples, posing more severe privacy risks compared to prior generative models such as GANs.
Figure 29: Illustration of data extraction attacks to LM agents.
Figure 30: Illustration of membership inference attacks (MIAs) to LM agents.
V-A2 Membership Inference Attacks (MIAs)